The price (in rands) in June of 1 kg packet of laundry in kwaDukuza is shown for each 5 years 2016 - 17.8, 2017 - 29.6, 2018 - 41.1, 2019 - 46.3, 2020 - 46.5 find the percentage increase in the price between each period


1
Expert's answer
2021-06-07T04:44:30-0400

Price of 1kg of laundry in kwaDukuza in 2016 is 17.8, in 2017 is 29.6, in 2018 is 41.1, in 2019 is 46.3 and in 2020 is 46.5.

So, the percentage increase in the price b/w each period is-:

In 2017 price percentage increase = "\dfrac{29.6}{17.8}\times 100 = 166.29"

so the price will be increase by 66.29 %.

In 2018 price percentage increase ="\dfrac{41.1}{29.6}\times 100 = 138.85"

so the price will be increase by 38.85 %.

In 2019 price percentage increase= "\dfrac{46.3}{41.1}\times 100 = 112.65"

so the price will be increase by 12.65%.

In 2020 price percentage increase= "\dfrac{46.5}{46.3}\times 100 = 100.43"

so the price will be increase by 0.43%.